OpenStudy (anonymous):

Find the common ratio of the arithmetic progression if the first term is 1/3 and the eighth term is 128/3.

OpenStudy (y2o2):

Tn = a + (n-1)d T1 = a = 1/3 T8 = (1/3) + 7d = (128/3) 7d = (127/3) d = (128/21)

OpenStudy (anonymous):

Thanks!